### v1

# Revisiting 360 Depth Estimation with PanoGabor: A New Fusion Perspective

Abstract
Depth estimation from a monocular 360 image is important to the perception of the entire 3D environment. However, the inherent distortion and large field of view (FoV) in 360 images pose great challenges for this task. To this end, existing mainstream solutions typically introduce additional perspective-based 360 representations (e.g., Cubemap) to achieve effective feature extraction. Nevertheless, regardless of the introduced representations, they eventually need to be unified into the equirectangular projection (ERP) format for the subsequent depth estimation, which inevitably reintroduces the troublesome distortions. In this work, we propose an oriented distortion-aware Gabor Fusion framework (PGFuse) to address the above challenges. First, we introduce Gabor filters that analyze texture in the frequency domain, thereby extending the receptive fields and enhancing depth cues. To address the reintroduced distortions, we design a linear latitude-aware distortion representation method to generate customized, distortion-aware Gabor filters (PanoGabor filters). Furthermore, we design a channel-wise and spatial-wise unidirectional fusion module (CS-UFM) that integrates the proposed PanoGabor filters to unify other representations into the ERP format, delivering effective and distortion-free features. Considering the orientation sensitivity of the Gabor transform, we introduce a spherical gradient constraint to stabilize this sensitivity. Experimental results on three popular indoor 360 benchmarks demonstrate the superiority of the proposed PGFuse to existing state-of-the-art solutions. Code can be available upon acceptance.
Index Terms:
Monocular, 360°, Depth Estimation, Distortion, Gabor Transform
Figure 1: Comparative of our proposed framework (shown in red) against other existing methods. Our method consistently demonstrates lower values across all error metrics. The compact red area highlights that our approach maintains competitive scores in computational efficiency (FLOPs), demonstrating a well-balanced trade-off between performance and computational cost.
I Introduction
360 images provide panoramic information of the surroundings with a single shot, while monocular 360 depth estimation enables the perception of the complete 3D scene information. However, mainstream models are specifically designed for perspective images and cannot address the challenges of distortions and the large FoV posed by 360 images due to fixed sampling positions and limited receptive fields. Consequently, addressing panoramic distortion and extending the receptive field have become primary concerns of current methodologies.
To mitigate the negative effects of distortion, BiFuse [1] introduced a dual-projection fusion scheme, which captures distortion-free features using an additional cubemap representation. However, its bidirectional fusion module imposes significant computational and memory burdens. To address this issue, Unifuse [2] proposes a more efficient unidirectional fusion module based on residual learning to reduce computational demands. Moreover, HRDfuse [3] replaces the cubemap projection with a more advanced Tangent Patch projection format (TP), achieving an effective fusion of TP and ERP features through a spatial feature alignment module. On the other hand, some researchers focus on introducing distortion awareness to transformer architectures with large receptive fields [4, 5, 6]. Considering the non-negligible computational and memory cost inherent in transformers, Elite360D [7] proposes a bi-projection bi-attention fusion module that integrates ERP and icosahedron projection (ICOSAP) features, combining the strengths of fusion-based schemes and transformer-based architectures. However, this type of fusion-based method, whether introducing new representations or proposing new fusion modules, overlooks the fact that all representations need to be integrated into the ERP format for fusion, which inevitably reintroduces distortion. Therefore, we argue that explicitly addressing the troublesome panoramic distortion in the fusion process can benefit the subsequent depth estimation.
To this end, we present PGFuse, an oriented panoramic Gabor-based fusion framework for monocular omnidirectional depth estimation. PGFuse addresses reintroduced distortions when integrating other representations into the ERP format. Benefiting from the frequency domain transformation, Gabor filters can extend the receptive field. Meanwhile, due to the strength of texture analysis capability, it can reduce the interference of surface texture, thereby enhancing the geometric depth cues. Besides, the tunable parameters of Gabor filters enable the creation of task-specific filters without additional parameters. However, the inherent distortion in 360 images poses a significant challenge for Gabor filter applications. To address this challenge, we design distortion-aware Gabor filters using a linear latitude-aware distortion representation method, as detailed in Eq. 2 and Algorithm 1, which adjusts the parameters of Gabor filters to achieve distortion awareness. These finely tuned Gabor filters, termed PanoGabor filters, are then incorporated as a convolutional layer. PanoGabor is embedded into a specially designed Channel-wise and Spatial-wise Unidirectional Fusion Module (CS-UFM) to mitigate the reintroduced distortions. More importantly, the parameters of PanoGabor can be optimized via dataset-driven training, ensuring seamless integration. Furthermore, we introduce a spherical gradient technique to stabilize the orientation sensitivity of the Gabor filters. Leveraging Gabor-based distortion-aware design and a perspective-based feature extraction strategy, PGfuse achieves the best-balanced trade-off between performance and computational cost, as shown in Figure. 1.
Extensive experiments evaluated on three popular indoor datasets—Stanford2D-3D [8], Matterport3D [9], and Structured3D [10]—demonstrate that our proposed scheme outperforms other state-of-the-art approaches. Our contributions can be summarized as follows.
  • We introduce PGFuse, a panoramic Gabor-based fusion framework that effectively addresses the reintroduced distortion problems that are often overlooked by previous methods.
  • To address the challenges of large FoV and panoramic distortion of 360 images, we design panorama-specific Gabor filters to analyze panoramic texture in the frequency domain, thereby extending the receptive fields and enhancing depth cues.
  • We design a channel-wise and spatial-wise unidirectional fusion module for the integration of other representations into ERP format, delivering effective and distortion-free features for the subsequent depth map estimation.
  • In popular benchmarks, our solution outperforms current state-of-the-art approaches.
Figure 2: Overview of the proposed framework.
II Related work
II-A Monocular 360 Depth Estimation
Panoramic distortion leads to notable performance degradation in perspective image-based depth estimation approaches. To this end, Zioulis et al. [11] highlighted the need to learn within the 360 domain, introducing the 3D60 dataset. Moreover, to mitigate distortion, researchers have adjusted the sampling positions of convolutional kernels for distortion awareness [12, 13, 14]. Some indoor monocular panoramic depth estimation methods [15, 16, 17] have developed various distortion-aware techniques to manage panoramic distortions, but most of these techniques are computational inefficient. Recently, advanced Transformer-based architectures [18] have broadened the receptive field, enabling models to extract features from 360 images. However, the traditional self-attention mechanism struggles to handle the complex distortion distribution in 360 images [4]. Enhanced distortion-aware transformer architectures [4, 5, 6, 19], while effective, often come with high computational and memory costs. To simplify models, some approaches [20, 21] advocate for estimating indoor depth from vertically compressed one-dimensional sequences, which leverages gravity alignment properties. However, this compression often fails to capture essential structural geometric information. To achieve a balanced trade-off between performance and efficiency, the mainstream approaches still prefer to develop a more efficient fusion-based framework. However, these fusion methods neglect the reintroduced distortion during the fusion process. Besides, although many of these approaches claim state-of-the-art performance, comparisons are often complicated by variations in datasets, preprocessing methods, and evaluation strategies, making fair comparisons challenging. To this end, Elite360D [7] provides a solid baseline. Therefore, in this paper, we follow the Elite360D method [7] to conduct the experiments.
II-B Gabor Filter in Image Processing
The Gabor filter is widely used in fields such as texture analysis, object detection, and face recognition, due to its robustness to varying angles and scales. Li et al. [22] suggested enhancing face recognition by learning the Gabor wavelet’s covariance matrix. In fabric defect detection, Chen et al. [23] introduced the Genetic Algorithm Gabor Faster R-CNN (Faster GG R-CNN), embedding Gabor kernels to reduce texture interference. Andrey et al. [24] proposed GaborNet, which uses Gabor filters to improve convergence and reduce the complexity of training in image recognition. To extend the receptive fields and enhance the geometric depth cues, we introduce Gabor filters to analyze texture in the frequency domain. Considering the special characteristics of 360 images, we propose a distortion-aware Gabor-based approach to address the challenging distortion issues.
III Methodology
In this section, we introduce the proposed PGFuse network. The complete framework, as illustrated in Figure 2, takes an ERP image as input and produces the corresponding predicted depth map. Our model follows a typical encoder-decoder scheme and consists of three main components: a feature extractor, a unidirectional fusion module, and an up-sampling decoder. To reduce the computational burden, we employ a shared-weight backbone to extract two sets of distortion-free cubemap features: one from the original perspective and the other horizontally rotated by 45°. These features are spatially aligned, unified into the ERP format, and then inputted into the CS-UMF module.
In the fusion module, we design PanoGabor filters to effectively handle the reintroduced distortion and enhance depth cues by analyzing textures in the frequency domain. Specifically, we introduce a linear latitude-aware distortion representation method that approximates the cosine value of latitude with a linear projection strategy. We then design an Algorithm 1 that uses this distortion representation to generate distortion-aware PanoGabor filters. These filters are further enclosed as a convolution layer and embedded in the CS-UFM module. Subsequently, the features processed by the CS-UFM module are fed into a multi-scale distortion-aware module (MS-DAM) [25] to integrate the shallow geometric features and the deep semantic features. These integrated features are then fed into the decoder, following a hierarchical architecture with multi-scale stages. Each stage consists of an upsampling layer and a convolution layer, followed by an ELU activation function. A final 3×3 convolution layer is used to produce the final depth map. Additionally, we project the ERP depth map onto a spherical surface and employ the Sobel operator to compute depth gradients in the nearest tangent patch, providing additional constraints and stabilizing the orientation sensitivity of the PanoGabor filters. The following sections provide detailed descriptions of the proposed modules.
Algorithm 1 Distortion-Aware PanoGabor Filters
Input: $c$: The defined distortion coefficient
Output: Convolution parameters
1:  Initialize convolution layer with given parameters
2:  Initialize Gabor parameters:$\theta\leftarrow\{\frac{\pi\cdot i}{8}\}_{i=1}^{8}$; $\psi=\pi\epsilon$; $f\leftarrow(\frac{\pi}{2})\cdot(\sqrt{2})^{\epsilon}\cdot(1+c)$; $\sigma\leftarrow\frac{\pi}{f+0.1}$
3:  Create meshgrid for $y$ and $x$: $[y,x]\leftarrow\text{meshgrid}((-1,0,1),(-1,0,1))$
4:  Reshape $\theta$, $\sigma$, $f$, and $\psi$ to match the required shape
5:  Calculate rotated $x$ and $y$: $x_{rot}\leftarrow x\cdot\cos(\theta)+y\cdot\sin(\theta)$; $y_{rot}\leftarrow-x\cdot\sin(\theta)+y\cdot\cos(\theta)$
6:  Calculate Gabor filter $g$:$g\leftarrow\exp\left(-0.5\cdot\left(\frac{x_{rot}^{2}+y_{rot}^{2}}{\sigma^{2}}\right)\right)\cdot\cos(f\cdot x_{rot}+\psi)$$g\leftarrow\frac{g}{2\cdot\pi\cdot\sigma^{2}}$
7:  Set convolutional parameters $\leftarrow g$
8:  return  Initialized distortion-aware Gabor-based convolutional parameters
III-A Distortion-aware PanoGabor Filters (PGF)
We aim to utilize the advanced texture analysis capabilities and inherent frequency domain transformation properties of Gabor filters to address the challenges posed by the large FoV in 360 images. However, due to the severe distortion, directly incorporating the standard Gabor filters into the learning frameworks for monocular 360 depth estimation would lead to unsatisfactory prediction. In contrast, the learnable Gabor filters can dynamically adapt to the distortions but deliver suboptimal performance because the distortion distribution pattern is hard to model in a data-driven way. Inspired by the spherical projection process, we propose to calculate the distortion coefficients for distortion-aware Gabor filter generation. Here, we propose the latitude-aware distortion representation method, which approximates the distortion coefficients as a function of the latitude.
The approximation derives from the equirectangular projection and the geometric properties of a sphere. Due to the spherical representation, the distance covered by a degree of longitude decreases from the equator towards the two poles. For a given latitude $\phi$, the distortion coefficient $c$ can be approximated by the reciprocal of the cosine of the latitude (the proof can be found in the suppl. material):
$c\approx\cos(\phi),$ (1)
the scale factor mirrors the increasing distortion in equirectangular projections towards the poles. However, in practice, we find that this nonlinear cosine representation introduces complex frequency changes making it more difficult to converge stably during training. Although it complies with certain distortion patterns, the instability in the optimization process may offset this advantage. To this end, we introduce a linear distortion representation method:
$c=|\frac{\phi\cdot\pi}{2}|,$ (2)
where $\phi\in[-\frac{\pi}{2},\frac{\pi}{2}]$. As detailed in Algorithm 1, we employ the defined distortion coefficient to control the generation of distortion-aware PanoGabor filters, thereby creating a bank of PanoGabor filters:
$PGF_{i}=$$\{g(x,y,f,\theta_{i},\psi,\sigma)\}_{i=1}^{8},$ (3)
where $g(x,y,f,\theta_{i},\psi,\sigma)$ indicates the Gabor transformation (preliminary definition can be found in the suppl. materials).
Figure 3: Visualization of the features extracted through the vanilla CNN layer and our proposed PanoGabor layer.
PanoGabor Convolution Layer. We further enclose the PanoGabor filters as an independent convolution layer. Given a set of input features $\mathcal{F}\in\mathbb{R}^{B\times C\times H\times W}$, we initially apply a standard convolutional layer to transform these features into channel-height dimension-matched features, denoted as $\mathcal{F^{\prime}}\in\mathbb{R}^{B\times H\times H\times W}$. For each channel, we then calculate a distortion coefficient based on the latitudinal distribution (i.e., $\phi\in[-\frac{\pi}{2},\frac{\pi}{2}]$). Using Algorithm 1, we generate a set of PanoGabor filters for each channel to compensate for the geometric distortion of spherical features. By further making the PanoGabor layer learnable, we enable the PanoGabor filters to become both distortion-aware and data-driven, thereby seamlessly integrating into the subsequent CS-UFM module for joint optimization. As illustrated in Figure 3, the proposed PanoGabor layer demonstrates its capability to reduce the interference of surface texture, thereby effectively extracting ERP features with explicit geometric structures from 360 images.
III-B Spherical Gradient Constraint
We have introduced our PanoGabor approach, which extends the original Gabor filters to be learnable and panorama-specific to achieve the dynamic perception of panoramic features. However, because of the sensitivity to the orientations, PanoGabor does not work well when the original gradient constraints are naively applied. Therefore, we perform a spherical gradient constraint for PanoGabor to assist in the dynamic parameter adjustment process. Specifically, we calculate these gradients on a pixel-level tangent plane. Mathematically, given a tangent sampling [26, 4, 19] function $f_{s}(*)$ over the sphere which can be employed to sample a 3×3 patch for each pixel, the spherical gradients can be defined as follows:
$G_{x}(p)$ $=$ $f_{s}(p)\times S_{x}$ (4)
$G_{y}(p)$ $=$ $f_{s}(p)\times S_{y},$ (5)
where $\phi$ and $\lambda$ represent latitude and longitude of a sphere, respectively; $S_{x},S_{y}$ represent the horizontal/vertical part of the Sobel operator.
III-C Channel-wise and Spatial-wise Unidirectional Fusion Module (CS-UFM)
The fusion module is designed to integrate two sets of ERP features, addressing reintroduced distortion, inconsistencies in cubemap boundaries, and pixel loss issues. This module first concatenates the two groups of input ERP features and uses a residual block to achieve an initial refinement. In particular, the block comprises a 1x1 convolution that compresses the channels to match the features’ height, facilitating the application of the PanoGabor layer. Following the 1×1 convolution layer, a 3x3 PanoGabor layer is used to process the features, aiming for distortion awareness and depth cues enhancement.
Subsequently, the enhanced features are multiplied by the mean values of the residual features on a spatial-wise basis to ensure a distortion-aware feature combination. To further refine the feature representation, a Squeeze-and-Excitation (SE) layer is applied, adaptively recalibrating channel-wise responses to emphasize important features and suppress redundant ones. The final stage involves a 1x1 convolution that produces the output feature map, effectively integrating the strengths of spatial and channel-wise processing.
This design enables the fusion module to enhance the analysis and integration of ERP features, making it particularly effective for subsequent depth estimation. The incorporation of PanoGabor filters and SE mechanisms provides a robust approach to enhance the module’s representational capacity, both channel-wise and spatial-wise.
Figure 4: The proposed CS-UFM module.
TABLE I: Quantitative comparison with the SOTA methods, including EGFormer [5], PanoFormer [4], BiFuse [1], BiFuse++ [27], UniFuse [2], OmniFusion [19], HRDFuse [28], and Elite360D [7].
Datasets Backbone Method Pub’Year Params (M) FLOPs $(\bar{G})$ Abs Rel $\downarrow$ Sq Rel $\downarrow$ $\mathrm{RMSE}\downarrow$ $\delta_{1}(\%)\uparrow$ $\delta_{2}(\%)\uparrow$ $\delta_{3}(\%)\uparrow$
Transformer EGFormer ICCV’23 15.39 66.21 0.1473 0.1517 0.6025 81.58 93.90 97.35
PanoFormer ECCV’22 20.38 81.09 0.1051 0.0966 0.4929 89.08 96.23 98.31
BiFuse CVPR’20 56.01 199.58 0.1126 0.0992 0.5027 88.00 96.13 98.47
BiFuse++ TPAMI’22 52.49 87.48 0.1123 0.0915 0.4853 88.12 96.56 98.69
M3D ResNet-34 UniFuse RAL’21 50.48 96.52 0.1144 0.0936 0.4835 87.85 96.59 98.73
OmniFusion CVPR’ 22 42.46 142.29 0.1161 0.1007 0.4931 87.72 96.15 98.44
HRDFuse CVPR’23 46.31 80.87 0.1172 0.0971 0.5025 86.74 96.17 98.49
Elite360D CVPR’24 25.54 65.29 0.1115 0.0914 0.4875 88.15 96.46 98.74
Ours - 30.75 75.67 0.1067 0.0858 0.4624 89.06 96.75 98.72
Transformer EGFormer ICCV’23 15.39 66.21 0.1528 0.1408 0.4974 81.85 93.38 97.36
PanoFormer ECCV’22 20.38 81.09 0.1122 0.0786 0.3945 88.74 95.84 98.59
OmniFusion CVPR’22 42.46 142.29 0.1154 0.0775 0.3809 86.74 96.03 98.71
S2D3D ResNet-34 UniFuse RAL’21 50.48 96.52 0.1124 0.0709 0.3555 87.06 97.04 98.99
Elite360D CVPR’24 25.51 65.28 0.1182 0.0728 0.3756 88.72 96.84 98.92
Ours - 30.75 75.67 0.1029 0.0599 0.3228 89.13 96.87 99.06
Transformer EGFormer ICCV’23 15.39 66.21 0.2205 0.4509 0.6841 79.79 90.71 94.55
PanoFormer ECCV’22 20.38 81.09 0.2549 0.4949 0.7937 74.70 89.15 93.97
Struct3D BiFuse CVPR’ 20 56.01 199.58 0.1573 0.2455 0.5213 85.91 94.00 96.72
ResNet-34 UniFuse RAL’ 21 50.48 96.52 0.1506 0.2319 0.5016 85.42 93.99 96.76
Elite360D CVPR’24 25.51 65.28 0.1480 0.2215 0.4961 87.41 94.34 96.66
Ours - 30.75 75.67 0.1401 0.2499 0.4394 87.78 94.49 96.82
III-D Loss Function
We employ the widely used reverse Huber Loss, and the proposed Spherical Gradient Loss to form our objective function. The standard loss functions of the former can be written as follows:
$\mathcal{L}_{\beta}(d,\hat{d})=\frac{1}{N}\sum\limits_{i\in N}\left\{\begin{array}[]{ll}|d_{i}-\hat{d_{i}}|&|d_{i}-\hat{d_{i}}|\leq\theta\\ \frac{(d_{i}-\hat{d_{i}})^{2}+\delta^{2}}{2\delta}&\text{ otherwise },\end{array}\right.$ (8)
where $d$/$\hat{d}$ is the ground-truth/predicted-value, $\theta$=0.2. For the Spherical Gradient Loss, we represent it as follows:
[…]
Ultimately, we formulate the objective function as follows:
[…]
where $\eta=0.5$; $G$: ground truth; $D$: predicted depth map.
IV Experiment
To validate the effectiveness of the proposed scheme, we conduct experiments on three popular datasets. The basic experiment settings are as follows.
We conduct all experiments on a single NVIDIA GeForce RTX 3090 GPU, implementing our approach using the PyTorch framework. We use the Adam optimizer [29] with an initialized learning rate of $1\times 10^{-4}$ to train our model. We keep the default settings for other hyper-parameters of the optimizer. During the training period, we use a batch size of 8 and employ left-right flipping, panoramic horizontal rotation, and luminance change for data augmentation, as previously proposed in works [2, 4]. The backbones we employed are all pre-trained on ImageNet-1K.
Datasets. For experimental validation, two real-world datasets (Stanford2D3D [8], and Matterport3D [9]) and a synthetic dataset (Structured3D [10]). We adopt a resolution of 1024×512 (width×height, 3 channels) of the single input panorama for all the datasets.
IV-A Evaluation Metrics
For evaluation, we follow the previous works [2, 4, 5, 7] and employ the standard evaluation metrics, including absolute relative error (Abs Rel), squared relative error (Sq Rel), Root Mean Squared Error (RMSE), and $\delta_{t},t\in\{1.25,1.25^{2},1.25^{3}\}$. Among them, $\delta_{t}$ provides a better reflection of the accuracy of the value per depth. RMSE indicates the average difference between the predicted and the actual value. Following Elite360D [7], we also report the number of parameters and FLOPS to evaluate the efficiency of our approach.
Figure 5: Qualitative comparison results. We exhibit the predicted depth maps of the Unifuse [2] and ours. All depth maps adopt a unified visualization method.
TABLE II: Ablation study.
Datasets Index Fusion Module Gabor Type Spherical Gradient $\mathrm{RMSE}\downarrow$ Sq Rel $\downarrow$ Abs Rel $\downarrow$ $\delta_{1}(\%)\uparrow$ $\delta_{2}(\%)\uparrow$ $\delta_{3}(\%)\uparrow$
a Cat 0.5021 0.0981 0.1125 87.29 95.72 98.19
b CS-UFM 0.4906 0.0971 0.1107 88.06 96.09 98.39
Mattterport3D c CS-UFM PanoGabor 0.4767 0.0903 0.1086 88.51 96.34 98.54
d CS-UFM PanoGabor 0.4624 0.0858 0.1067 89.06 96.75 98.72
e CS-UFM Normal Gabor 0.4848 0.0917 0.1089 88.25 96.25 98.46
IV-B Comparison Analysis
Quantitative Results. We exhibit the quantitative comparison results in Table I, which demonstrate that our model outperforms all other state-of-the-art solutions on most metrics. Our PGFuse method significantly outperforms compared methods. Specifically, on the real-world Matterport3D and Stanford2D3D datasets, our approach demonstrates notable improvements in the important RMSE and $\delta_{1}$ metrics. Additionally, it achieves comparable results on the virtual Structured3D dataset, showcasing its robustness and effectiveness across various environments. These results highlight the efficacy of PGFuse in monocular 360-depth estimation across diverse and challenging datasets. Although our performance on the loosest accuracy metric $\delta_{3}$ is slightly lower than the Elite360D [7] and the Unifuse [2], it performs better than the two methods on the other two tighter accuracy metrics ($\delta_{1}$ and $\delta_{2}$). Note that in the small-scale realistic Standford2D3D dataset, PGFuse outperforms other methods in all metrics. In [7], the authors argue that due to the limited data in the Stanford2D3D data set, the transformer-based approach can hardly provide accurate global perception, while the core components of our method are CNN-based and cannot be influenced by this negative effect. As for the Structred3D dataset, we can observe that our approach performs badly on the Sq Rel metric than the Unifuse and Elite360D, we suspect it might be related to the pre-processing operation. For a fair comparison, we follow the Elite360D to process the data (normalize the data and scale it to 10m), which confuses the model in estimating an accurate depth range.
[…]
Complexity. We show the complexity comparison results in Table I. We can observe that those transformer-based architectures may have fewer parameters and FLOPs. However, the complexity of these approaches is quadric to the input pixels, which limits its application to high-resolution images. In addition, their graphics memory cost is huge, which makes it difficult to use in practice. The other SoTA approach, Bifuse, has the largest parameters and huge FLOPs. And the training time is unbearable for large-scale datasets. Although our scheme has slightly more parameters and FLOPs than the Elite360D, the proposed approach performs better than Elite360D, especially on the RMSE metric, by a significant margin. In addition, we argue that the core fusion component of the Elite360D is attention-based, similar to the transformer-based frameworks, resulting in fewer parameters and FLOPs.
Figure 6: Visualization of the heatmap of the features before/after the CS-UFM module. Because each face in the cubemap projection is processed individually, resulting in discontinuous ERP features. The CS-UFM module can address this issue effectively.
IV-C Ablation Study
To demonstrate the effectiveness of the main proposed components, we follow Unifuse [2] and Elite360D [7] to design the ablation study on the Matterport3D dataset [9]. All the models are with the same conditions and trained to converge. As illustrated in Figure 2, we remove the fusion stage from the full framework to serve as the baseline for conducting the ablation study. The experimental results of the baseline are presented as ‘a’ in Table II. We further add the proposed CS-UFM module, a version without the PanoGabor layer embedded as ‘b’. Index ‘c’ refers to the use of the full CS-UFM module. Index ‘d’ indicates that we employ the proposed spherical gradient constraint to stabilize the orientation sensitivity of the Gabor filters. Additionally, in ‘e’, all PanoGabor layers are replaced with normal Gabor filters to further verify the effectiveness of the designed PanoGarbor.
Effectiveness of the CS-UFM module. From Table II, we can observe that the CS-UFM module can obtain an improvement even without the designed PanoGabor layer. This is because the fusion module can effectively fuse the two sets of features due to the channel-wise and spatial-wise designs. Moreover, when further embedding the proposed PanoGabor layer, our approach gets a significant improvement across all the evaluation metrics, verifying the necessity of addressing the reintroduced distortion problems during the fusion process.
Effectiveness of the PanoGabor layer. Comparing the b and e in Table II, we can observe that benefiting from the frequency domain transformation along with the capability of reducing surface texture interference, the normal Gabor layer can benefit the fusion process. From d and e, we can observe that our designed PanoGabor performs better than the normal ones due to the distortion-aware designs.
Effectiveness of the spherical gradient constraint. Recently, some works [4, 7] prefer to employ a gradient-based constraint to sharpen the depth edges. This strategy aims to extract the horizontal/vertical structures to design the objective function. However, since the Gabor filters are orientation-sensitive, we design the spherical gradient constraint to incorporate the panoramic geometry to stabilize this orientation sensitivity. From Table II, we can observe that the PanoGabor with the proposed spherical gradient constraint performs much better than that without.
Various Backbones. Following Elite360D [7], we also adopt various backbones for our framework. From Table III, we can observe that our scheme can benefit from employing more powerful backbones.
Trade-off between performance and efficiency. From Table IV, Elite360D has obtained a significant improvement by employing EfficientNet-B5 as a feature extractor. However, it brings unbearable memory costs. In contrast, our approach can get a superior performance just by replacing a group of input features without additional memory cost. This also demonstrates the robustness and scalability of our PGFuse.
TABLE III: Ablation study on various backbones.
Datasets Backbones Abs Rel $\downarrow$ Sq Rel $\downarrow$ $\mathrm{RMSE}\downarrow$ $\delta_{1}(\%)\uparrow$
ResNet-18 0.1125 0.0977 0.4933 87.78
M3D ResNet-34 0.1067 0.0858 0.4624 89.06
ResNet-50 0.1058 0.0841 0.4618 88.93
TABLE IV: More advanced performance comparison results. * represent that Elite360D employs Efficient-B5 as backbone. indicates that we replace one of the cubemap features with extracted ERP features.
Datasets Method Abs Rel $\downarrow$ Sq Rel $\downarrow$ $\mathrm{RMSE}\downarrow$ $\delta_{1}(\%)\uparrow$
M3D Elite360D* 0.1048 0.0805 0.4524 89.92
M3D Ours 0.0972 0.0801 0.4480 90.67
V Conclusion
In this paper, we propose a Gabor-based framework, termed PGFuse for 360 depth estimation. Since the existing fuse-based scheme ignores the reintroduced distortion when integrating other representations into ERP format, our proposed PGFuse embeds a Gabor-based layer to a unidirectional fusion module to explicitly deal with the reintroduced distortion. Specifically, to make the Gabor filters adapt to 360 images, we first introduce a linear latitude-aware representation to model the panoramic distortion distribution pattern. Then we leverage this representation to generate customized Gabor filters. These tuned PanoGabor can enhance the depth cues in the frequency domain and achieve effective distortion awareness. Therefore, our fusion module can deliver spatial-continuous and content-consistent ERP features. Extensive experiment results demonstrate that our approach outperforms current SoTA approaches.